### v1

# Zero-Shot Video Editing through Adaptive Sliding Score Distillation

Abstract
The burgeoning field of text-based video generation (T2V) has reignited significant interest in the research of controllable video editing. Although pre-trained T2V-based editing models have achieved efficient editing capabilities, current works are still plagued by two major challenges. Firstly, the inherent limitations of T2V models lead to content inconsistencies and motion discontinuities between frames. Secondly, the notorious issue of over-editing significantly disrupts areas that are intended to remain unaltered. To address these challenges, our work aims to explore a robust video-based editing paradigm based on score distillation. Specifically, we propose an Adaptive Sliding Score Distillation strategy, which not only enhances the stability of T2V supervision but also incorporates both global and local video guidance to mitigate the impact of generation errors. Additionally, we modify the self-attention layers during the editing process to further preserve the key features of the original video. Extensive experiments demonstrate that these strategies enable us to effectively address the aforementioned challenges, achieving superior editing performance compared to existing state-of-the-art methods. For more information, please visit our project homepage: https://nips24videoedit.github.io/zeroshot_videoedit/
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
1 Introduction
Video editing, as a task of interactive and controllable generation, has consistently garnered significant attention from the Artificial Intelligence Generated Content (AIGC) community. These techniques enable the modification of existing visual content to meet specific expectations, thereby finding extensive applications in fields such as photography, cinematography, and artistic creation. Despite substantial early explorations, the recent proliferation of Text-to-Image (T2I) and Text-to-Video (T2V) diffusion models has once again spurred a significant amount of related research.
Building on the foundation of these pre-trained diffusion models, existing research can be broadly categorized into Zero-Shot and One-Shot based on their training paradigms. Zero-shot approaches [22, 16, 9] utilize DDIM (Denoising Diffusion Implict Model) Inversion [11] to retain partial information from the original video and incorporate editing guidance as well as temporal priors through attention manipulation and inter-frame attention operations without any training  [5, 14, 31]. In contrast, One-shot approaches [34, 32, 24] attempt to fine-tune pre-trained models to learn subject-independent motion patterns, thereby enabling the editing of video content through prompt replacement. Despite the efficiency and impressive performance of Gaussian noise inversion in diffusion models, this paradigm also introduces two interrelated issues in video editing. Firstly, despite attempts to incorporate DDIM inversion, noise-based inference inevitably disrupts the content and temporal continuity of the source video; Secondly, due to the inherent issues in open-source video generation models, the generated videos frequently exhibit content misalignment and motion flicker, which are even more pronounced to existing video editing works.
A naive solution is to design a more robust update paradigm to replace noise-based ones, which is structure-agnostic and can be seamlessly integrated into most existing works. As in [21] in text-to-3D works, Score Distillation Loss (SDS) is a widely adopted method that aligns with our objective of maintaining consistency. Furthermore, a similar concept Delta Denoising Score (DDS) has also been proven effective in image editing [10], where DDS achieves image editing by directly optimizing the parameters in the source image, ensuring that -edited regions remain largely unaffected. Similarly, this work develops a novel video score distillation paradigm to achieve video-based parameters updating directly. Despite the remarkable success of the score distillation in 3D generation and image editing, experiments have shown that directly applying them [10, 21] with existing open-source text-to-video models [30, 27] yields unsatisfactory results. The edited videos exhibit spotty or blocky bright artifacts, which we refer to as Contaminations. We hypothesize that this issue stems from the weak robustness and unstable generative performance in T2V models, in contrast to T2I models, which adversely affects gradient computation. Undeniably, this issue is exacerbated in video-based update paradigms, as the latent codes of the original video, being at a high resolution and low abstraction level closer to the pixel layer, are particularly sensitive to these errors. Furthermore, image score distillation methods lack video-specific priors and global understanding capabilities, thereby exacerbating the contaminations of the source information.
Figure 1: Our video editing results. More experimental results and complete videos can be found on our project homepage.
This work designs an Adaptive Sliding Score Distillation (ASSD) method, which aims to introduce more robust supervision and video priors to mitigate the aforementioned issues. Specifically, we firstly introduce more stable T2I models to assist in latent code optimization, thereby compensating for the instability of video generation models. To adapt SDS for video data, we propose a local score distillation strategy based on the sliding time-window, which incorporates inter-frame continuity priors during the update process. However, merely relying on local priors is often inefficient, as a significant amount of contaminations in complex video data is difficult to address. Therefore, we design a global gradient-based contaminations prediction strategy to guide the entire distillation process. Additionally, to ensure that the information in the source video is not adversely affected by erroneous generation from the T2V model, specific spatial self-attention layers in the denoising U-Net within the editing branch are modified. Through our proposed weighted attention fusion, we incorporate reference branch information into the gradient computation for optimizing the latent code, correcting errors in the editing branch, and thus enhancing the quality of the update gradients. Extensive experiments on both real and synthetic video editing demonstrate the effectiveness of our proposed method, achieving comparable or superior results to other video editing approaches. In summary, the contributions can be summarized as:
  • We innovatively propose a zero-shot pipeline that addresses video editing by directly optimizing the original video through score distillation.
  • To tackle the contaminations in T2V score distillation, we design ASSD (Adaptive Sliding Score Distillation) to introduce global and local priors for gradient calculation. Additionally, we employ the stable T2I model to provide auxiliary guidance during the editing process.
  • We modify certain spatial self-attention layers in the denoising U-Net, enhancing the robustness of the editing process by preserving information from the original video.
  • Extensive experiments on real-world videos demonstrate that our method achieves comparable or better performance than existing methods in terms of content consistency, motion continuity, adherence to the editing text, and overall video quality.
Figure 2: Overview of Pipeline. Our pipeline includes an upper reference branch and a lower editing branch. We use ASSD for T2V models and DDS for T2I models to jointly derive the gradients for updating the latent code of the video. Additionally, we modify the computation in the spatial self-attention layers of the editing branch to incorporate information from the reference branch.
2 Related Work
Text-to-Video Generation.
Early works on conditional video generation utilized frameworks such as variational autoencoders [15], generative adversarial networks [35, 4, 25, 28], or autoregressive transformers [7, 13, 33, 29], which were trained on small domain-specific datasets. Despite achieving promising results, these approaches struggled to generalize to scenarios beyond the training datasets. Recently, numerous efforts have been made to extend the success of diffusion models in text-to-image generation to the task of text-to-video generation [8, 20]. By training on large-scale image and video datasets, these models have achieved remarkable generation results and controllability over the generated content. To accommodate the input data type of videos, the prevailing methods involve extending the convolutional and attention layers of text-to-image diffusion models along the temporal dimension, such as incorporating temporal convolutions, temporal attention layers, or introducing attention mechanisms between different frames of the video [26, 6, 3, 30, 36]. In this work, our objective is to harness the capabilities of pre-trained text-to-video diffusion models to perform zero-shot editing of original videos, given the corresponding text prompts and edited text prompts. This is achieved solely by utilizing the outputs of the pre-trained model, without subjecting it to any further training.
Text-based Video Editing.
With the success achieved in the field of text-based video generation, many works have attempted to utilize text-based image and video generation models for text-conditioned video editing. Given a reference video with corresponding text prompts and a target text prompt, the video editing task aims to edit the reference video according to the target text prompt.
Tune-A-Video [34] introduces the ability to model temporal information for 2D diffusion models by fine-tuning pre-trained text-to-image diffusion models on reference videos while using DDIM inversion to provide layout and motion guidance from the reference video for editing. Video-P2P [16] adapts the cross-attention control method from the image editing work Prompt-to-Prompt [11] to video editing, employing a dual-branch pipeline that combines the unedited region from the reconstruction branch and the edited region latent features from the editing branch according to the cross-attention map, achieving editing effects in the edited regions while preserving the unedited regions of the reference video. FateZero [22], on the other hand, directly manipulates the intermediate features during the diffusion model generation process under the guidance of the cross-attention map, realizing zero-shot video editing while preserving information from the reference video using DDIM inversion. FLATTEN [5] adopts a similar pipeline to FateZero but introduces guidance from optical flow predicted from the video. TokenFlow [9] and CoDeF [18] embody similar ideas: TokenFlow performs editing on key frames during the generation process and propagates the edits from the key frames to the entire video through feature matching. CoDeF decomposes the video into a canonical content field containing the static content of the entire video and a time-dependent deformation field that includes the transformations needed from the canonical to each frame of the video. By editing the canonical content field and utilizing the deformation field, the editing results are propagated to the entire video.
These works have achieved excellent video editing results, but they typically require executing a complete generation process to obtain the edited video. Although methods such as DDIM inversion have been employed, they still lose some information from the original video, resulting in the problem of maintaining unedited content and motion of the edited subject from the original video. In contrast to their work, our approach directly optimizes the original video by leveraging the guidance provided by pre-trained text-to-video models, better preserving the unedited content from the original video and utilizing the motion information of the edited subject from the original video to achieve stronger inter-frame motion continuity.
3 Method
3.1 Preliminaries
Text-conditioned Latent Video Diffusion Models.
We utilize pre-trained text-conditioned latent video diffusion models [30, 27] to provide guidance for video editing. The text-conditioned latent video diffusion model performs forward noise addition and reverse denoising operations in the latent space of an autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$. We denote a video sequence consisting of $N$ frames as $\bm{X}_{0}=\{\bm{x}_{0}^{1},...,\bm{x}_{0}^{N}\}$. During the training phase, the latent input $\bm{Z}_{0}=\{\bm{z}_{0}^{1},...,\bm{z}_{0}^{N}\}=\mathcal{E}(\bm{X}_{0})$ is perturbed to $\bm{Z}_{t}=\{\bm{z}_{t}^{1},...,\bm{z}_{t}^{N}\}$ through the forward diffusion process:
[…]
for $t\in[1,T]$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}1-\beta_{s}$, and $\beta_{s}$ is the variance schedule for timestep $s$. Subsequently, the 3D denoising U-Net is trained to predict the added noise given condition text embedding $y$ using the following objective function:
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(\bm{X}_{0}),y,\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),t}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)\|_{2}^{2}\right],$ (2)
where $\bm{\epsilon}_{\theta}$ represents the 3D denoising U-Net. In text-conditioned generative inference, CFG (Classifier-Free Guidance) [12] is commonly used to strengthen the effect of the condition. The noise prediction in this context is calculated using the following formula:
$\bm{\epsilon}_{\theta}^{w}(\bm{Z}_{t},t,y)=\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)+w(\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)),$ (3)
where $\emptyset$ is null text embedding and $w$ is guidance scale.
Score Distillation for Image Editing.
The SDS loss was first introduced in the text-to-3D work DreamFusion [21], utilizing a pre-trained text-to-image diffusion model to provide update guidance for implicit 3D representations such as NeRF (Neural Radiance Fields) [17]. However, recent work [10] demonstrates that this score distillation loss is equally effective for image editing. Denote the pre-trained T2I model as $\bm{\epsilon}_{\theta}$. Given the parameterized latent code $\bm{z}_{0}(\phi)$ of an image to be edited, the source text embedding $y$, and the target text embedding $y^{*}$, we can use the SDS loss to directly optimize the image to be edited, which can be expressed as:
$\mathcal{L}_{\mathrm{SDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}\right\|_{2}^{2},\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),$ (4)
for $t\in[1,T]$. Although the SDS loss provides update guidance for $\bm{z}_{0}(\phi)$ to align with the target text prompt, it often results in over-saturation and blurriness, making the image appear cartoon-like [10]. To address this issue, DDS introduces an additional reference branch aligned with the source text prompt. By calculating the delta score between the model outputs of the two branches, it provides more accurate guidance for updates in the edited parts of $\bm{z}_{0}(\phi)$, while preserving the unedited parts. Let the reference latent code be denoted as $\bm{\hat{z}}_{0}$, which is initialized as the original latent code of the image. The DDS loss can be formulated as:
$\mathcal{L}_{\mathrm{DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{z}}_{t},t,y)\right\|_{2}^{2}.$ (5)
Figure 3: Adaptive Sliding Score Distillation. ASSD adaptively extracts masks for regions with significant gradient changes based on the temporal variation of the update gradients. It then employs a sliding window-based smoothing operation to mitigate errors in the noise predictions from the reference branch.
3.2 Adaptive Sliding Score Distillation
Using a pre-trained text-conditioned latent video diffusion model, we can naturally extend the DDS loss used for image editing to videos. In this work, we use the open-source T2V diffusion models ModelScopeT2V and ZeroScope, denoted as $\bm{\epsilon}_{\theta}$. The video DDS loss can be expressed as:
$\mathcal{L}_{\mathrm{Video-DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}\left(\bm{Z}_{t}(\phi),t,y^{*}\right)-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{Z}}_{t},t,y)\right\|_{2}^{2},$ (6)
and the gradient used to update $\bm{Z}_{0}(\phi)$ can be calculated as:
$\nabla_{\phi}\mathcal{L}_{\mathrm{Video-DDS}}=\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.$ (7)
The calculation of this gradient involves a one-step noise prediction from the pre-trained T2V diffusion model. Due to the significantly weaker performance of existing open-source T2V diffusion models in video generation tasks compared to the performance of Stable Diffusion in image generation, this one-step noise prediction contains substantial errors, and becomes particularly pronounced at large timesteps. This leads to unacceptable outcomes when directly transferring the DDS loss to video editing, as evidenced by the results shown in the ablation study in Fig. 5. In the editing branch, the significant variation of the noise predictions at different locations is because the predicted noise also contains information related to the edits. In contrast, the noise predictions in the reference branch should not exhibit such drastic fluctuations. Therefore, we employ a sliding window approach to smooth the noise prediction results of the reference branch along the temporal dimension. The width of the sliding window should not be set too large, as excessive smoothing could lead to blurring of the video. Thus, we selected a small sliding window width of 3. For convenience, we temporarily omit the subscript $\theta$ and superscript $\omega$, let $\bm{\hat{\epsilon}}_{t}^{i}$ denote the noise prediction for the $i$-th frame of the video by the reference branch at timestep $t$. The sliding window-based smoothing operation for this frame is then defined as:
$\operatorname{SS}(\bm{\hat{\epsilon}}_{t}^{i})=\gamma_{i-1}\cdot\bm{\hat{\epsilon}}_{t}^{i-1}+\gamma_{i}\cdot\bm{\hat{\epsilon}}_{t}^{i}+\gamma_{i+1}\cdot\bm{\hat{\epsilon}}_{t}^{i+1},$ (8)
where $\gamma_{i-1},\gamma_{i},\gamma_{i+1}$ are hyper-parameters that control the degree of smoothing, and $\operatorname{SS}(\cdot)$ denotes our sliding window-based smoothing operation. However, applying this sliding window-based smoothing across the entire video can still introduce unnecessary blurring, especially in regions without large prediction errors. To mitigate the impact, guidance is needed to determine where to apply the sliding window-based smoothing. Considering that these prediction errors are typically concentrated around the edited parts, where the update gradients exhibit significant variations along the temporal dimension, we can construct a mask to guide the range of the sliding window-based smoothing operation accordingly. We adaptively obtain the mask through the following process:
$\bm{m}=\left\{\begin{array}[]{ll}1,&\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)<\mathrm{max}\left(\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)\right)-\tau\\ 0,&\text{otherwise}\end{array},\right.$ (9)
where $\bm{\hat{\epsilon}}_{t}$ is the noise prediction for all frames output by the reference branch, $\bm{{\epsilon}}_{t}$ is the noise prediction for all frames output by the editing branch, $\operatorname{max}(\cdot)$ denotes the maximization operation along the temporal dimension, and $\tau$ is a hyper-parameter used to control the size of the mask’s effective range. When $\tau$ is smaller, the mask’s effective range is smaller and more stringent, and vice versa. Based on this adaptively extracted mask, the final gradient used to update $\bm{Z}_{0}(\phi)$ is represented as follows:
$\begin{split}\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}&=\left(\bm{m}\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\mathrm{SS}\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\right.\\ &\quad+\left.(\bm{1}-\bm{m})\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.\end{split}$ (10)
In our proposed adaptive sliding score distillation (ASSD), we mitigate the impact of model prediction errors using a sliding window along the temporal dimension. Additionally, we construct a mask adaptively to minimize the negative effects of this smoothing operation, such as blurring in regions of the video that do not require smoothing. This approach effectively reduces the occurrence of contaminations in the edited video results.
3.3 Weighted Attention Fusion
Although the proposed ASSD effectively reduces errors in the update gradient, local editing errors and changes in unedited parts still occur, which can be observed in Fig. 5. It is desirable to preserve some information in the original video, including the layout of the original video, the details in the unedited parts, and some attributes in the edited parts such as the motion pattern of the edited object. Moreover, some features from the original video can also be leveraged to correct erroneous parts in the edited video to further enhance the editing quality. To better preserve the information and leverage some features in the original video, we introduce a weighted attention fusion.
In the spatial attention layers of ModelScopeT2V or Zeroscope, the self-attention for the latent code $\bm{z}^{i}$ corresponding to the $i$-th frame in the video is calculated as:
$\mathrm{SelfAttn}_{i}=\mathrm{Softmax}(\frac{\bm{W}_{q}(\bm{z}^{i})\bm{W}_{k}(\bm{z}^{i})^{\top}}{\sqrt{d}})\bm{W}_{v}(\bm{z}^{i}),$ (11)
where $\bm{W}_{q}(\cdot)$, $\bm{W}_{k}(\cdot)$, $\bm{W}_{v}(\cdot)$ are linear projections, and $d$ is the scaling factor. We use $\bm{z}^{i}$ to denote the latent code corresponding to the $i$-th frame in the editing branch of ASSD. Correspondingly, we use $\bm{\hat{z}}^{i}$ to denote the latent code corresponding to the $i$-th frame in the reference branch of ASSD. The proposed reference-attention between them is defined as:
$\mathrm{RefAttn}_{i}=\mathrm{Softmax}(\frac{\bm{W}_{q}(\bm{z}^{i})\bm{W}_{k}(\bm{\hat{z}}^{i})^{\top}}{\sqrt{d}})\bm{W}_{v}(\bm{\hat{z}}^{i}).$ (12)
Then, our weighted attention fusion can be implemented as:
$\mathrm{AttnFusion}_{i}=\lambda\cdot\mathrm{SelfAttn}_{i}+(1-\lambda)\cdot\mathrm{RefAttn}_{i},$ (13)
where $\lambda\in[0,1]$. This module allows for the fusion of features from the editing branch with those from the reference branch. We replaced the self-attention calculations in the last few spatial attention layers of the U-Net in the editing branch with our proposed weighted attention fusion, and find this design significantly improves the quality of the edited videos.
3.4 Joint Guidance Incorporating Stable Diffusion
As validated in DDS, the one-step predicted noise provided by Stable Diffusion can offer sufficient quality gradient information for image editing. Therefore, we propose a joint guidance strategy that introduces Stable Diffusion to provide additional gradients for updating $\bm{Z}_{0}(\phi)$. Stable Diffusion uses DDS to separately calculate the gradient for updating each frame in the video. For the $i$-th frame in the video, the gradient can be calculated as follows,
$\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}=\left(\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{z}_{t}^{i},y^{*},t\right)-\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{\hat{z}}_{t}^{i},y,t\right)\right)\frac{\partial{\bm{z}^{i}}}{\partial\phi},$ (14)
where $\bm{\epsilon}_{\Omega}$ denotes a Stable Diffusion V1.5 in our implementation. Finally, the total gradient used to update $\bm{Z}_{0}(\phi)$ can be expressed as:
$\nabla_{\phi}\mathcal{L}_{\mathrm{total}}=\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}+\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}.$ (15)
Figure 4: Qualitative Comparison. Comparison of our ASSD with five baseline methods. Our method achieves a good balance between editing and preserving unedited information in the original video. Additionally, it possesses a certain capability for geometric shape editing.
4 Experiments
4.1 Experiments Setting
We collected approximately 30 text-video pairs from public datasets such as DAVIS [19], WebVid [2], and YouTube [1], and extracted 8 or 16 evenly spaced frames from each video for our evaluation data. We used ModelscopeT2V and Zeroscope as our base models. For optimizing the latent code of a video, we employed the SGD optimizer with a learning rate of 0.1, running the optimization for 200 steps, the process takes approximately 3 to 4 minutes on a single V100 GPU.
We evaluated our method alongside five high-performing and representative video editing methods. Tune-A-Video (TAV) [34] fine-tunes the denoising U-Net on a single video to learn the motion patterns of the video to be edited. FateZero [22] achieves zero-shot video editing through DDIM Inversion and attention manipulation techniques. FLATTEN [5] utilizes an external optical flow detection model to enable more accurate attention manipulation. TokenFlow [9] ensures the consistency of the edited video by modifying keyframes and propagating these modifications throughout the entire video. ControlVideo [37] utilizes information such as depth and edges from the original video to control the editing results.
4.2 Qualitative Experiment
Fig. 4 presents the qualitative comparison results between our method and other SOTA baselines. Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and preservation of unedited regions.
Table 1: Quantitative comparisons of our ASSD against SOTA video edit methods.
Methods Auto Metrics User Study
CLIP-Text Align CLIP-Consistency Alignment Consistency Preservation
TAV [34] 0.2988 0.9086 19.06 2.64 4.25
FateZero [22] 0.2738 0.9324 20.63 26.17 30.62
ControlVideo [37] 0.2916 0.9301 8.48 5.30 3.70
FLATTEN [5] 0.2931 0.9249 14.82 17.99 11.11
TokenFlow [9] 0.2910 0.9561 12.67 20.11 24.40
ASSD (Ours) 0.2899 0.9113 24.34 27.79 25.93
4.3 Quantitative Experiment
Auto Metrics.
We utilized the CLIP model [23] for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score. CLIP Text Alignment Score is defined as the mean cosine similarity between the CLIP image embeddings of the frames in the edited video and the CLIP text embedding of the target text prompt. This metric measures how well the edited video aligns with the target text prompt. CLIP Frame Consistency Score is defined as the mean cosine similarity between the CLIP image embeddings of frames within the video. This metric assesses the inter-frame consistency of the edited video. The results of the automatic metric are reported in Tab. 1.
User Study.
Due to the current lack of appropriate evaluation metrics in the field of video generation and editing, automatic assessments based on CLIP often do not align well with the perceptual quality of videos. Therefore, we primarily validate the effectiveness of our method through user studies. We invited 40 volunteers to participate in our user study. The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment with the target text prompt. (2) Inter-frame consistency of the edited video. (3) Preservation of the original video’s motion patterns and unedited regions. For each criterion, the volunteers chose the best-performing edited result from those produced by our method and the five baselines. We then calculated the proportion of times each method’s edited results were selected as the best for each criterion. As shown in Tab. 1, our method achieves comparable or superior performance across all three evaluation metrics compared to other SOTA methods.
Figure 5: Abalation of ASSD, Weighted Attention Fusion and Joint Guidance with Stable Diffusion. In most cases, using all the components we propose achieves the best balance between editing and minimizing contaminations.
4.4 Ablation Study
We conduct ablation studies below to validate the effect of each proposed component. As shown in the first row of experimental results in Fig. 5, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video. The results in the second row show that by using ASSD instead of DDS, and smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality. When we further incorporate the proposed weighted attention fusion mechanism, using features from the reference branch, the contaminations are further reduced. Additionally, we observe that the edited video retains more information from the original video, such as colors and other details. Due to the significantly weaker robustness and generalization to different text conditions of current open-source text-to-video models compared to Stable Diffusion, the latter can sometimes provide higher-quality update gradients for certain text prompts, despite lacking the ability to model temporal information. As shown in the last row of Fig. 5, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects.
5 Conclusion and Limitations
We propose a new score distillation-based paradigm for video editing and introduce a novel score distillation method called Adaptive Sliding Score Distillation (ASSD). ASSD mitigates errors in the gradient information provided by text-to-video models through an adaptive sliding window-based smoothing operation and adaptively extracted masks. Our pipeline also incorporates auxiliary guidance from Stable Diffusion to provide more precise gradients for updating the latent code of original video. Additionally, to better capture the information from the original video to correct errors in the edited video and further improve the quality of the update gradients, we employ proposed weighted attention fusion. This modifies certain spatial self-attention layers in pre-trained text-to-video models to support the fusion of information between the editing branch and the reference branch. Experimental results validate the effectiveness and potential of our new paradigm in video editing tasks, as well as the efficacy of our proposed components.
Limitations
It should be noted that the performance of our method heavily relies on the capabilities of the text-to-video models providing the gradient information for editing. The effectiveness of such methods relies on the premise that the pre-trained text-to-video model can provide an update gradient that, despite containing errors, is still effective. Given the current lack of sufficiently powerful open-source text-to-video models, the upper bound of our method’s performance is significantly constrained.

### v2

@@
-Zero-Shot Video Editing through Adaptive Sliding Score Distillation
+Adaptive Sliding Score Distillation for Zero-Shot Video Editing
 Abstract
-The burgeoning field of text-based video generation (T2V) has reignited significant interest in the research of controllable video editing.
-Although pre-trained T2V-based editing models have achieved efficient editing capabilities, current works are still plagued by two major challenges.
-Firstly, the inherent limitations of T2V models lead to content inconsistencies and motion discontinuities between frames.
-Secondly, the notorious issue of over-editing significantly disrupts areas that are intended to remain unaltered.
-To address these challenges, our work aims to explore a robust video-based editing paradigm based on score distillation.
-Specifically, we propose an Adaptive Sliding Score Distillation strategy, which not only enhances the stability of T2V supervision but also incorporates both global and local video guidance to mitigate the impact of generation errors.
-Additionally, we modify the self-attention layers during the editing process to further preserve the key features of the original video.
-Extensive experiments demonstrate that these strategies enable us to effectively address the aforementioned challenges, achieving superior editing performance compared to existing state-of-the-art methods.
-For more information, please visit our project homepage: https://nips24videoedit.github.io/zeroshot_videoedit/
-11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
-1 Introduction
-Video editing, as a task of interactive and controllable generation, has consistently garnered significant attention from the Artificial Intelligence Generated Content (AIGC) community.
-These techniques enable the modification of existing visual content to meet specific expectations, thereby finding extensive applications in fields such as photography, cinematography, and artistic creation.
-Despite substantial early explorations, the recent proliferation of Text-to-Image (T2I) and Text-to-Video (T2V) diffusion models has once again spurred a significant amount of related research.
-Building on the foundation of these pre-trained diffusion models, existing research can be broadly categorized into Zero-Shot and One-Shot based on their training paradigms.
-Zero-shot approaches [22, 16, 9] utilize DDIM (Denoising Diffusion Implict Model) Inversion [11] to retain partial information from the original video and incorporate editing guidance as well as temporal priors through attention manipulation and inter-frame attention operations without any training  [5, 14, 31].
-In contrast, One-shot approaches [34, 32, 24] attempt to fine-tune pre-trained models to learn subject-independent motion patterns, thereby enabling the editing of video content through prompt replacement.
-Despite the efficiency and impressive performance of Gaussian noise inversion in diffusion models, this paradigm also introduces two interrelated issues in video editing.
-Firstly, despite attempts to incorporate DDIM inversion, noise-based inference inevitably disrupts the content and temporal continuity of the source video; Secondly, due to the inherent issues in open-source video generation models, the generated videos frequently exhibit content misalignment and motion flicker, which are even more pronounced to existing video editing works.
-A naive solution is to design a more robust update paradigm to replace noise-based ones, which is structure-agnostic and can be seamlessly integrated into most existing works.
-As in [21] in text-to-3D works, Score Distillation Loss (SDS) is a widely adopted method that aligns with our objective of maintaining consistency.
-Furthermore, a similar concept Delta Denoising Score (DDS) has also been proven effective in image editing [10], where DDS achieves image editing by directly optimizing the parameters in the source image, ensuring that -edited regions remain largely unaffected.
-Similarly, this work develops a novel video score distillation paradigm to achieve video-based parameters updating directly.
-Despite the remarkable success of the score distillation in 3D generation and image editing, experiments have shown that directly applying them [10, 21] with existing open-source text-to-video models [30, 27] yields unsatisfactory results.
-The edited videos exhibit spotty or blocky bright artifacts, which we refer to as Contaminations.
-We hypothesize that this issue stems from the weak robustness and unstable generative performance in T2V models, in contrast to T2I models, which adversely affects gradient computation.
-Undeniably, this issue is exacerbated in video-based update paradigms, as the latent codes of the original video, being at a high resolution and low abstraction level closer to the pixel layer, are particularly sensitive to these errors.
-Furthermore, image score distillation methods lack video-specific priors and global understanding capabilities, thereby exacerbating the contaminations of the source information.
-Figure 1: Our video editing results. More experimental results and complete videos can be found on our project homepage.
-This work designs an Adaptive Sliding Score Distillation (ASSD) method, which aims to introduce more robust supervision and video priors to mitigate the aforementioned issues.
-Specifically, we firstly introduce more stable T2I models to assist in latent code optimization, thereby compensating for the instability of video generation models.
-To adapt SDS for video data, we propose a local score distillation strategy based on the sliding time-window, which incorporates inter-frame continuity priors during the update process.
-However, merely relying on local priors is often inefficient, as a significant amount of contaminations in complex video data is difficult to address.
-Therefore, we design a global gradient-based contaminations prediction strategy to guide the entire distillation process.
-Additionally, to ensure that the information in the source video is not adversely affected by erroneous generation from the T2V model, specific spatial self-attention layers in the denoising U-Net within the editing branch are modified.
-Through our proposed weighted attention fusion, we incorporate reference branch information into the gradient computation for optimizing the latent code, correcting errors in the editing branch, and thus enhancing the quality of the update gradients.
-Extensive experiments on both real and synthetic video editing demonstrate the effectiveness of our proposed method, achieving comparable or superior results to other video editing approaches.
-In summary, the contributions can be summarized as:
-  • We innovatively propose a zero-shot pipeline that addresses video editing by directly optimizing the original video through score distillation.
-  • To tackle the contaminations in T2V score distillation, we design ASSD (Adaptive Sliding Score Distillation) to introduce global and local priors for gradient calculation.
-Additionally, we employ the stable T2I model to provide auxiliary guidance during the editing process.
-  • We modify certain spatial self-attention layers in the denoising U-Net, enhancing the robustness of the editing process by preserving information from the original video.
-  • Extensive experiments on real-world videos demonstrate that our method achieves comparable or better performance than existing methods in terms of content consistency, motion continuity, adherence to the editing text, and overall video quality.
-Figure 2: Overview of Pipeline.
-Our pipeline includes an upper reference branch and a lower editing branch.
-We use ASSD for T2V models and DDS for T2I models to jointly derive the gradients for updating the latent code of the video.
-Additionally, we modify the computation in the spatial self-attention layers of the editing branch to incorporate information from the reference branch.
-2 Related Work
-Text-to-Video Generation.
-Early works on conditional video generation utilized frameworks such as variational autoencoders [15], generative adversarial networks [35, 4, 25, 28], or autoregressive transformers [7, 13, 33, 29], which were trained on small domain-specific datasets.
-Despite achieving promising results, these approaches struggled to generalize to scenarios beyond the training datasets.
-Recently, numerous efforts have been made to extend the success of diffusion models in text-to-image generation to the task of text-to-video generation [8, 20].
-By training on large-scale image and video datasets, these models have achieved remarkable generation results and controllability over the generated content.
-To accommodate the input data type of videos, the prevailing methods involve extending the convolutional and attention layers of text-to-image diffusion models along the temporal dimension, such as incorporating temporal convolutions, temporal attention layers, or introducing attention mechanisms between different frames of the video [26, 6, 3, 30, 36].
-In this work, our objective is to harness the capabilities of pre-trained text-to-video diffusion models to perform zero-shot editing of original videos, given the corresponding text prompts and edited text prompts.
-This is achieved solely by utilizing the outputs of the pre-trained model, without subjecting it to any further training.
+The rapidly evolving field of Text-to-Video generation (T2V) has catalyzed renewed interest in controllable video editing research.
+While the application of editing prompts to guide diffusion model denoising has gained prominence, mirroring advancements in image editing, this noise-based inference process inherently compromises the original video’s integrity, resulting in unintended over-editing and temporal discontinuities.
+To address these challenges, this study proposes a novel paradigm of video-based score distillation, facilitating direct manipulation of original video content.
+Specifically, distinguishing it from image-based score distillation, we propose an Adaptive Sliding Score Distillation strategy, which incorporates both global and local video guidance to reduce the impact of editing errors.
+Combined with our proposed Image-based Joint Guidance mechanism, it has the ability to mitigate the inherent instability of the T2V model and single-step sampling.
+Additionally, we design a Weighted Attention Fusion module to further preserve the key features of the original video and avoid over-editing.
+Extensive experiments demonstrate that these strategies effectively address existing challenges, achieving superior performance compared to current state-of-the-art methods.
+Figure 1: Our Video Editing Results. More experimental results and complete videos can be found in supplementary material.
+Introduction
+Video editing, an interactive and controllable generation task, has consistently attracted significant attention from the Artificial Intelligence Generated Content (AIGC) community.
+These techniques allow the modification of visual content to meet specific expectations, finding extensive applications in photography, cinematography, and artistic creation.
+The recent proliferation of Text-to-Image (T2I) and Text-to-Video (T2V) diffusion models has significantly propelled this field.
+Building on these pre-trained models, current research primarily falls into two categories: Zero-Shot and One-Shot approaches.
+Zero-shot methods (Qi et al.
+2023; Liu et al.
+2023; Geyer et al.
+2023) utilize DDIM Inversion (Hertz et al.
+2022) to preserve original video information, incorporating editing guidance and inter-frame attention during inference for continuous video editing without additional training (Cong et al.
+2023; Jeong and Ye 2023; Wang et al.
+2023b).
+In contrast, One-shot methods (Wu et al.
+2023; Wei et al.
+2023; Ren et al.
+2024) fine-tune pre-trained models to learn subject-independent motion patterns, enabling video content editing through prompt replacement.
+A common thread among these approaches is their reliance on the inference process of diffusion models, which aims to controllably recover edited images from the noise—a process termed the noise-based paradigm.
+Despite employing DDIM inversion-related techniques, these methods intrinsically result in degradation of the original input structure.
+Furthermore, cumulative uncertainty during inference also exacerbates this effect, consequently inducing the notorious issues of over-editing and discontinuity in video editing.
+A naive solution is to design a robust update paradigm, which possesses the intuitive capacity for direct video-based (or latent code) optimization by predicting update-gradients while preserving consistency in the non-edited areas.
+Inspired by text-to-3D works (Poole et al.
+2022), Score Distillation Sampling (SDS) is a widely adopted method for maintaining consistency.
+Furthermore, a similar concept Delta Denoising Score (DDS) has also been proven effective in image editing (Hertz, Aberman, and Cohen-Or 2023), where DDS achieves image editing by directly optimizing the parameters in the source image, ensuring that unedited regions remain largely unaffected.
+However, the direct transposition of this paradigm to video editing presents significant challenges.
+Extensive empirical investigation reveals that edited videos often exhibit over-editing and pronounced artifacts, a phenomenon we term Contamination in Fig.
+4.
+This issue likely stems from the weak robustness and unstable performance of T2V diffusion models, affecting gradient computation.
+Moreover, the SDS paradigm, utilizing single-step sampling in diffusion models, inherently exacerbates this problem.
+Undeniably, these characteristics are critical in video-based update paradigms, as the the original video structures are more fragile, making them particularly sensitive to these errors.
+Fig.
+5 empirically validates this hypothesis, showing that T2V diffusion models like ModelscopeT2V and Zeroscope exhibit significant errors in one-step predicted noise used for video update-gradients at larger timesteps, compared to Stable Diffusion.
+This work introduces an Adaptive Sliding Score Distillation (ASSD) methodology, designed to incorporate more robust supervision and video priors to address the aforementioned challenges, as in Fig.
+1.
+Specifically, we propose a localized score distillation strategy utilizing sliding time-windows, which integrates inter-frame continuity priors during the update process to mitigate abrupt gradient fluctuations within local temporal sequences.
+However, recognizing that local priors alone often prove insufficient for complex video data, we develop a global gradient-based Contamination prediction strategy to guide the distillation process.
+Based on this methodology, we design an Image-based Joint Guidance mechanism, which enhances the upload-gradients from the selected frames perspective.
+Conversely, ASSD also mitigates the inherent discontinuities in multi-images editing in Joint Guidance mechanism.
+Moreover, to further mitigate potential adverse effects stemming from suboptimal T2V model and preserve the integrity of non-edited regions in the original video, we modify specific spatial self-attention layers within the denoising U-Net architecture in the editing branch.
+Through the implementation of a Weighted Attention Fusion module, we incorporate reference branch guidance into the updating-gradients computation in editing branch, thereby rectifying errors in the editing branch.
+Fig.
+6 illustrates the application of our method directly to the original video, demonstrating the editing process.
+Comprehensive experiments conduct on both real and synthetic video editing tasks substantiate the efficacy of our proposed method, yielding superior results compared to alternative approaches.
+In summary, our contributions are:
+  • We propose a novel video-based optimizing pipeline for zero-shot video editing through score distillation.
+  • To address T2V score distillation Contamination, we design Adaptive Sliding Score Distillation, introducing global and local priors for updating-gradients calculation, and design Image-based Joint Guidance mechanism for auxiliary guidance.
+  • We design a Weighted Attention Fusion module to enhance robustness and preserve original video structures.
+  • Extensive experiments show our method achieves superior performance to existing methods in content consistency, motion continuity, adherence to editing text, and overall video quality.
+Figure 2: Overview of Pipeline and proposed ASSD Loss.
+Our pipeline comprises reference and editing branches.
+We employ ASSD for T2V diffusion model and Image-based Joint Guidance to derive gradients for latent code updates, while some of the editing branch’s spatial self-attention layers are modified to incorporate reference branch information using Weighted Attention Fusion mechanism.
+The ASSD illustration on the right showcases sliding windows as semi-transparent orange rectangle, with $\mathbf{m}$ indicating areas of substantial gradient change, extracted adaptively.
+Related Work
 Text-based Video Editing.
-With the success achieved in the field of text-based video generation, many works have attempted to utilize text-based image and video generation models for text-conditioned video editing.
-Given a reference video with corresponding text prompts and a target text prompt, the video editing task aims to edit the reference video according to the target text prompt.
-Tune-A-Video [34] introduces the ability to model temporal information for 2D diffusion models by fine-tuning pre-trained text-to-image diffusion models on reference videos while using DDIM inversion to provide layout and motion guidance from the reference video for editing.
-Video-P2P [16] adapts the cross-attention control method from the image editing work Prompt-to-Prompt [11] to video editing, employing a dual-branch pipeline that combines the unedited region from the reconstruction branch and the edited region latent features from the editing branch according to the cross-attention map, achieving editing effects in the edited regions while preserving the unedited regions of the reference video.
-FateZero [22], on the other hand, directly manipulates the intermediate features during the diffusion model generation process under the guidance of the cross-attention map, realizing zero-shot video editing while preserving information from the reference video using DDIM inversion.
-FLATTEN [5] adopts a similar pipeline to FateZero but introduces guidance from optical flow predicted from the video.
-TokenFlow [9] and CoDeF [18] embody similar ideas: TokenFlow performs editing on key frames during the generation process and propagates the edits from the key frames to the entire video through feature matching.
-CoDeF decomposes the video into a canonical content field containing the static content of the entire video and a time-dependent deformation field that includes the transformations needed from the canonical to each frame of the video.
-By editing the canonical content field and utilizing the deformation field, the editing results are propagated to the entire video.
-These works have achieved excellent video editing results, but they typically require executing a complete generation process to obtain the edited video.
-Although methods such as DDIM inversion have been employed, they still lose some information from the original video, resulting in the problem of maintaining unedited content and motion of the edited subject from the original video.
-In contrast to their work, our approach directly optimizes the original video by leveraging the guidance provided by pre-trained text-to-video models, better preserving the unedited content from the original video and utilizing the motion information of the edited subject from the original video to achieve stronger inter-frame motion continuity.
-3 Method
-3.1 Preliminaries
+With the success achieved in the field of Text-based Video generation(Hong et al.
+2022; Rombach et al.
+2022; Saharia et al.
+2022; Wang et al.
+2023a; Pika 2023; Zhang et al.
+2023a), many works have attempted to utilize T2I and T2V diffusion models for text-conditioned video editing.
+Tune-A-Video (Wu et al.
+2023) pioneered one-shot video editing by fine-tuning pre-trained text-to-image diffusion models on reference videos, utilizing DDIM inversion for layout and motion guidance.
+Subsequent zero-shot methods emerged, employing DDIM Inversion-like techniques to preserve original video information while manipulating the inference process.
+Video-P2P (Liu et al.
+2023) adapted cross-attention control from Prompt-to-Prompt (Hertz et al.
+2022), using a dual-branch pipeline.
+FateZero (Qi et al.
+2023) manipulated intermediate features during diffusion model generation with cross-attention guidance.
+FLATTEN (Cong et al.
+2023) incorporated optical flow guidance.
+TokenFlow (Geyer et al.
+2023) propagated edits through feature matching, while CoDeF (Ouyang et al.
+2023) decomposed videos into static content and time-dependent deformation fields for editing and propagation.
+These noise-based methods achieve excellent video editing but require a complete noise addition and denoising process, often losing some original video information and failing to fully maintain unedited content and motion.
+Therefore, this work explores a video-based updating paradigm using guidance from pre-trained T2V diffusion models, better preserving unedited content and inter-frame motion continuity.
+Additional related works are discussed in the appendix.
+Preliminaries
 Text-conditioned Latent Video Diffusion Models.
-We utilize pre-trained text-conditioned latent video diffusion models [30, 27] to provide guidance for video editing.
-The text-conditioned latent video diffusion model performs forward noise addition and reverse denoising operations in the latent space of an autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$.
+We utilize pre-trained text-conditioned latent video diffusion models (Wang et al.
+2023a; Sterling 2023) to provide guidance for video editing.
+These models perform forward noise addition and reverse denoising operations in the latent space based on an autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$.
 We denote a video sequence consisting of $N$ frames as $\bm{X}_{0}=\{\bm{x}_{0}^{1},...,\bm{x}_{0}^{N}\}$.
@@
 for $t\in[1,T]$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}1-\beta_{s}$, and $\beta_{s}$ is the variance schedule for timestep $s$.
-Subsequently, the 3D denoising U-Net is trained to predict the added noise given condition text embedding $y$ using the following objective function:
+Subsequently, the denoising U-Net is trained to predict the added noise given condition text embedding $y$ using the following objective function:
 $\mathcal{L}=\mathbb{E}_{\mathcal{E}(\bm{X}_{0}),y,\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),t}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)\|_{2}^{2}\right],$ (2)
 where $\bm{\epsilon}_{\theta}$ represents the 3D denoising U-Net.
-In text-conditioned generative inference, CFG (Classifier-Free Guidance) [12] is commonly used to strengthen the effect of the condition.
-The noise prediction in this context is calculated using the following formula:
-$\bm{\epsilon}_{\theta}^{w}(\bm{Z}_{t},t,y)=\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)+w(\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)),$ (3)
-where $\emptyset$ is null text embedding and $w$ is guidance scale.
 Score Distillation for Image Editing.
-The SDS loss was first introduced in the text-to-3D work DreamFusion [21], utilizing a pre-trained text-to-image diffusion model to provide update guidance for implicit 3D representations such as NeRF (Neural Radiance Fields) [17].
-However, recent work [10] demonstrates that this score distillation loss is equally effective for image editing.
-Denote the pre-trained T2I model as $\bm{\epsilon}_{\theta}$.
-Given the parameterized latent code $\bm{z}_{0}(\phi)$ of an image to be edited, the source text embedding $y$, and the target text embedding $y^{*}$, we can use the SDS loss to directly optimize the image to be edited, which can be expressed as:
-$\mathcal{L}_{\mathrm{SDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}\right\|_{2}^{2},\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),$ (4)
-for $t\in[1,T]$.
-Although the SDS loss provides update guidance for $\bm{z}_{0}(\phi)$ to align with the target text prompt, it often results in over-saturation and blurriness, making the image appear cartoon-like [10].
+The SDS loss was first introduced in the text-to-3D work DreamFusion (Poole et al.
+2022), utilizing a pre-trained text-to-image diffusion model to provide update guidance for implicit 3D representations such as NeRF (Neural Radiance Fields) (Mildenhall et al.
+2021).
+Recent work (Hertz, Aberman, and Cohen-Or 2023) has demonstrated the efficacy of score distillation loss in image editing tasks.
+However, directly applying the SDS loss to compute gradient information for image editing leads to over-saturated and blurred results, imparting a cartoon-like appearance to the images.
 To address this issue, DDS introduces an additional reference branch aligned with the source text prompt.
+Denote the pre-trained T2I diffusion model as $\bm{\epsilon}_{\theta}$.
+Given the parameterized latent code $\bm{z}_{0}(\phi)$ of an image to be edited, the source text embedding $y$, and the target text embedding $y^{*}$.
 By calculating the delta score between the model outputs of the two branches, it provides more accurate guidance for updates in the edited parts of $\bm{z}_{0}(\phi)$, while preserving the unedited parts.
@@
 $\mathcal{L}_{\mathrm{DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{z}}_{t},t,y)\right\|_{2}^{2}.$ (3)
-Figure 3: Adaptive Sliding Score Distillation.
-ASSD adaptively extracts masks for regions with significant gradient changes based on the temporal variation of the update gradients.
-It then employs a sliding window-based smoothing operation to mitigate errors in the noise predictions from the reference branch.
-3.2 Adaptive Sliding Score Distillation
-Using a pre-trained text-conditioned latent video diffusion model, we can naturally extend the DDS loss used for image editing to videos.
-In this work, we use the open-source T2V diffusion models ModelScopeT2V and ZeroScope, denoted as $\bm{\epsilon}_{\theta}$.
-The video DDS loss can be expressed as:
+for $t\in[1,T]$, and $w$ represents the classifier-free guidance weight (Ho and Salimans 2022) utilized in the sampling process of the diffusion model.
+Method
+Overview
+Given an original video $\bm{X}_{0}$ with its corresponding text prompt $y$ and a target text prompt $y^{*}$, our method directly edits the latent code $\bm{Z}_{0}$ of the original video to align with the target text prompt.
+As illustrated in Fig.
+2, ASSD and Image-based Joint Guidance provide more stable update gradients for video editing, mitigating over-editing.
+Weighted Attention Fusion incorporates features from the reference branch, thus better preserving the original video information.
+Adaptive Sliding Score Distillation
+To introduce SDS-related update paradigms in video-based editing, a natural approach is to extend DDS to temporal inputs.
+Based on the advanced open-source T2V diffusion models ModelScopeT2V and ZeroScope, denoted as $\bm{\epsilon}_{\theta}$, video DDS loss can be expressed as:
 $\mathcal{L}_{\mathrm{Video-DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}\left(\bm{Z}_{t}(\phi),t,y^{*}\right)-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{Z}}_{t},t,y)\right\|_{2}^{2},$ (4)
@@
 $\nabla_{\phi}\mathcal{L}_{\mathrm{Video-DDS}}=\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.$ (5)
-The calculation of this gradient involves a one-step noise prediction from the pre-trained T2V diffusion model.
-Due to the significantly weaker performance of existing open-source T2V diffusion models in video generation tasks compared to the performance of Stable Diffusion in image generation, this one-step noise prediction contains substantial errors, and becomes particularly pronounced at large timesteps.
-This leads to unacceptable outcomes when directly transferring the DDS loss to video editing, as evidenced by the results shown in the ablation study in Fig.
-5.
-In the editing branch, the significant variation of the noise predictions at different locations is because the predicted noise also contains information related to the edits.
-In contrast, the noise predictions in the reference branch should not exhibit such drastic fluctuations.
-Therefore, we employ a sliding window approach to smooth the noise prediction results of the reference branch along the temporal dimension.
-The width of the sliding window should not be set too large, as excessive smoothing could lead to blurring of the video.
-Thus, we selected a small sliding window width of 3.
-For convenience, we temporarily omit the subscript $\theta$ and superscript $\omega$, let $\bm{\hat{\epsilon}}_{t}^{i}$ denote the noise prediction for the $i$-th frame of the video by the reference branch at timestep $t$.
-The sliding window-based smoothing operation for this frame is then defined as:
-$\operatorname{SS}(\bm{\hat{\epsilon}}_{t}^{i})=\gamma_{i-1}\cdot\bm{\hat{\epsilon}}_{t}^{i-1}+\gamma_{i}\cdot\bm{\hat{\epsilon}}_{t}^{i}+\gamma_{i+1}\cdot\bm{\hat{\epsilon}}_{t}^{i+1},$ (8)
-where $\gamma_{i-1},\gamma_{i},\gamma_{i+1}$ are hyper-parameters that control the degree of smoothing, and $\operatorname{SS}(\cdot)$ denotes our sliding window-based smoothing operation.
-However, applying this sliding window-based smoothing across the entire video can still introduce unnecessary blurring, especially in regions without large prediction errors.
-To mitigate the impact, guidance is needed to determine where to apply the sliding window-based smoothing.
-Considering that these prediction errors are typically concentrated around the edited parts, where the update gradients exhibit significant variations along the temporal dimension, we can construct a mask to guide the range of the sliding window-based smoothing operation accordingly.
-We adaptively obtain the mask through the following process:
+Nevertheless, in practical applications, while this paradigm demonstrates the ability to preserve non-edited regions, as illustrated in the first row of Fig.
+4, it consistently encounters challenges such as over-editing (manifested as unintended changes in vehicle color) and the emergence of undesirable artifacts proximal to the edited object.
+The observed anomalies can be ascribed to a confluence of factors.
+Primarily, open-source T2V diffusion models demonstrate markedly inferior stability in video generation tasks relative to their Stable Diffusion counterparts in image generation.
+Moreover, the single-step sampling strategy inherent to SDS-analogous paradigms serves to amplify this instability, resulting in a compounded degradation of output fidelity.
+While the integration of frame-by-frame image editing techniques and multi-step sampling strategies offers a promising avenue for improvement, these methodologies are constrained by challenges in maintaining inter-frame coherence and computational efficiency.
+Consequently, it is imperative to explore a novel video score distillation strategy, namely Adaptive Sliding Score Distillation, which demonstrates potential in alleviating the aforementioned coupled challenges.
+Specifically, we analyze diverse branches, including editing and reference branches, to examine the predictive noise distribution.
+In the editing branch, noise predictions exhibit significant temporal variability due to the incorporation of edit-related guidance.
+Conversely, the reference branch’s predictions are expected to demonstrate substantially less temporal heterogeneity.
+Therefore, we employ a sliding time-window to smooth the reference branch’s noise prediction, which reflects the inherent continuity of the reference video and influences the smoothness of upload gradients.
+The width of the sliding window should not be set too large, as excessive smoothing could lead to over-blurring of the video.
+For simplicity, we temporarily omit the subscript $\theta$ and superscript $\omega$.
+Let $\bm{\hat{\epsilon}}_{t}^{i}$ denote the noise prediction for the $i$-th frame in the reference branch at timestep $t$.
+The sliding time-window smoothing operation for the current frame,
+$\begin{split}\operatorname{Smooth}(\bm{\hat{\epsilon}}_{t}^{i})=\gamma_{i-1}\cdot\bm{\hat{\epsilon}}_{t}^{i-1}+\gamma_{i}\cdot\bm{\hat{\epsilon}}_{t}^{i}&+\gamma_{i+1}\cdot\bm{\hat{\epsilon}}_{t}^{i+1},\\ \gamma_{i-1}+\gamma_{i}+\gamma_{i+1}&=1,\end{split}$ (6)
+where $\gamma_{i-1},\gamma_{i},\gamma_{i+1}$ are hyper-parameters that control the degree of smoothing, and $\operatorname{Smooth}(\cdot)$ denotes our sliding time-window smoothing operation.
+In our experiments, $\gamma_{i}$ is set to $0.67$, while $\gamma_{i-1}$ and $\gamma_{i+1}$ are equal.
+However, applying this window-based smoothing strategy across the entire video may introduce unnecessary blurring, as both the regions intended for editing and those erroneously edited theoretically produce discontinuous update gradients in the temporal dimension, while the former is necessary in the editing task.
+This is challenging because these two types of regions are often intertwined.
+To mitigate this effect, an additional strategy for determining smooth regions is indispensable.
+Intuitively, regions with undesired over-editing and artifacts often exhibit larger update gradients compared to correctly edited areas.
+Therefore, we attempt to filter out these regions by applying an adaptive mask, $m$, which incorporates global video guidance and has the capability to handle diverse editing objectives.
+Specifically,
 $\bm{m}=\left\{\begin{array}[]{ll}1,&\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)<\mathrm{max}\left(\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)\right)-\tau\\ 0,&\text{otherwise}\end{array},\right.$ (7)
@@
 Based on this adaptively extracted mask, the final gradient used to update $\bm{Z}_{0}(\phi)$ is represented as follows:
-$\begin{split}\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}&=\left(\bm{m}\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\mathrm{SS}\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\right.\\ &\quad+\left.(\bm{1}-\bm{m})\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.\end{split}$ (10)
-In our proposed adaptive sliding score distillation (ASSD), we mitigate the impact of model prediction errors using a sliding window along the temporal dimension.
-Additionally, we construct a mask adaptively to minimize the negative effects of this smoothing operation, such as blurring in regions of the video that do not require smoothing.
-This approach effectively reduces the occurrence of contaminations in the edited video results.
-3.3 Weighted Attention Fusion
-Although the proposed ASSD effectively reduces errors in the update gradient, local editing errors and changes in unedited parts still occur, which can be observed in Fig.
-5.
+$\begin{split}\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}=\left(\bm{m}\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\mathrm{Smooth}\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\right.\\ \quad+\left.(\bm{1}-\bm{m})\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.\end{split}$ (8)
+In Adaptive Sliding Score Distillation (ASSD), we mitigate the impact of upload gradient errors by leveraging local time-windows information.
+In addition, we adaptively construct a mask based on the global guidance to minimize the negative effects of the over-smoothing.
+This method effectively reduces contaminations in the edited video results.
+Figure 3: Qualitative Comparison.
+Our method achieves a good balance between editing and preserving unedited information in the original video.
+Additionally, it possesses a certain capability for geometric shape editing.
+Image-based Joint Guidance
+As validated in DDS, the one-step predicted noise provided by Stable Diffusion can offer sufficient quality gradient information for image editing.
+Therefore, we propose a joint guidance strategy that introduces Stable Diffusion to provide more stable gradients for updating $\bm{Z}_{0}(\phi)$, which uses DDS to separately calculate the gradient for updating selected frames in the video.
+In our experiments, we select keyframes using a fixed step size to maintain efficiency.
+For the $i$-th selected frame in the video, the gradient is,
+$\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}=\left(\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{z}_{t}^{i},y^{*},t\right)-\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{\hat{z}}_{t}^{i},y,t\right)\right)\frac{\partial{\bm{z}^{i}}}{\partial\phi},$ (9)
+where $\bm{\epsilon}_{\Omega}$ denotes a Stable Diffusion V1.5 in our implementation. Finally, the total gradient used to update $\bm{Z}_{0}(\phi)$ can be expressed as:
+$\nabla_{\phi}\mathcal{L}_{\mathrm{total}}=\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}+\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}.$ (10)
+Weighted Attention Fusion
+Although the proposed ASSD effectively reduces errors in the update gradients, local editing errors and over-editing still occur, which can be observed in Fig.
+4.
 It is desirable to preserve some information in the original video, including the layout of the original video, the details in the unedited parts, and some attributes in the edited parts such as the motion pattern of the edited object.
-Moreover, some features from the original video can also be leveraged to correct erroneous parts in the edited video to further enhance the editing quality.
-To better preserve the information and leverage some features in the original video, we introduce a weighted attention fusion.
-In the spatial attention layers of ModelScopeT2V or Zeroscope, the self-attention for the latent code $\bm{z}^{i}$ corresponding to the $i$-th frame in the video is calculated as:
+Intuitively, features from the original video can be utilized to reduce over-editing, thereby further enhancing the editing quality.
+To better preserve the information and leverage features in the original video, we introduce a Weighted Attention Fusion module.
+Specifically, in the spatial attention layers of ModelScopeT2V or Zeroscope, the self-attention for the latent code $\bm{z}^{i}$ corresponding to the $i$-th frame in the video is:
 $\mathrm{SelfAttn}_{i}=\mathrm{Softmax}(\frac{\bm{W}_{q}(\bm{z}^{i})\bm{W}_{k}(\bm{z}^{i})^{\top}}{\sqrt{d}})\bm{W}_{v}(\bm{z}^{i}),$ (11)
 where $\bm{W}_{q}(\cdot)$, $\bm{W}_{k}(\cdot)$, $\bm{W}_{v}(\cdot)$ are linear projections, and $d$ is the scaling factor.
-We use $\bm{z}^{i}$ to denote the latent code corresponding to the $i$-th frame in the editing branch of ASSD.
-Correspondingly, we use $\bm{\hat{z}}^{i}$ to denote the latent code corresponding to the $i$-th frame in the reference branch of ASSD.
+We use $\bm{z}^{i}$ to denote the latent code corresponding to the $i$-th frame in the editing branch of ASSD and use $\bm{\hat{z}}^{i}$ to denote the latent code corresponding to the $i$-th frame in the reference branch of ASSD.
 The proposed reference-attention between them is defined as:
@@
 Then, our weighted attention fusion can be implemented as:
-$\mathrm{AttnFusion}_{i}=\lambda\cdot\mathrm{SelfAttn}_{i}+(1-\lambda)\cdot\mathrm{RefAttn}_{i},$ (13)
-where $\lambda\in[0,1]$.
+$\mathrm{AttnFusion}_{i}=(1-\lambda)\cdot\mathrm{SelfAttn}_{i}+\lambda\cdot\mathrm{RefAttn}_{i},$ (13)
+where $\lambda\in[0,1]$, it controls the balance between the extent of editing and the preservation of information from the original video.
+In the majority of cases, we set it to 0.8.
 This module allows for the fusion of features from the editing branch with those from the reference branch.
-We replaced the self-attention calculations in the last few spatial attention layers of the U-Net in the editing branch with our proposed weighted attention fusion, and find this design significantly improves the quality of the edited videos.
-3.4 Joint Guidance Incorporating Stable Diffusion
-As validated in DDS, the one-step predicted noise provided by Stable Diffusion can offer sufficient quality gradient information for image editing.
-Therefore, we propose a joint guidance strategy that introduces Stable Diffusion to provide additional gradients for updating $\bm{Z}_{0}(\phi)$.
-Stable Diffusion uses DDS to separately calculate the gradient for updating each frame in the video.
-For the $i$-th frame in the video, the gradient can be calculated as follows,
-$\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}=\left(\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{z}_{t}^{i},y^{*},t\right)-\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{\hat{z}}_{t}^{i},y,t\right)\right)\frac{\partial{\bm{z}^{i}}}{\partial\phi},$ (14)
-where $\bm{\epsilon}_{\Omega}$ denotes a Stable Diffusion V1.5 in our implementation. Finally, the total gradient used to update $\bm{Z}_{0}(\phi)$ can be expressed as:
-$\nabla_{\phi}\mathcal{L}_{\mathrm{total}}=\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}+\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}.$ (15)
-Figure 4: Qualitative Comparison.
-Comparison of our ASSD with five baseline methods.
-Our method achieves a good balance between editing and preserving unedited information in the original video.
-Additionally, it possesses a certain capability for geometric shape editing.
-4 Experiments
-4.1 Experiments Setting
-We collected approximately 30 text-video pairs from public datasets such as DAVIS [19], WebVid [2], and YouTube [1], and extracted 8 or 16 evenly spaced frames from each video for our evaluation data.
-We used ModelscopeT2V and Zeroscope as our base models.
-For optimizing the latent code of a video, we employed the SGD optimizer with a learning rate of 0.1, running the optimization for 200 steps, the process takes approximately 3 to 4 minutes on a single V100 GPU.
-We evaluated our method alongside five high-performing and representative video editing methods.
-Tune-A-Video (TAV) [34] fine-tunes the denoising U-Net on a single video to learn the motion patterns of the video to be edited.
-FateZero [22] achieves zero-shot video editing through DDIM Inversion and attention manipulation techniques.
-FLATTEN [5] utilizes an external optical flow detection model to enable more accurate attention manipulation.
-TokenFlow [9] ensures the consistency of the edited video by modifying keyframes and propagating these modifications throughout the entire video.
-ControlVideo [37] utilizes information such as depth and edges from the original video to control the editing results.
-4.2 Qualitative Experiment
-Fig.
-4 presents the qualitative comparison results between our method and other SOTA baselines.
-Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and preservation of unedited regions.
+We replaced the self-attention calculations in the last few spatial attention layers of the U-Net only in the editing branch with our proposed weighted attention fusion, and find this design significantly improves the quality of the edited videos.
+Experiments
 Table 1: Quantitative comparisons of our ASSD against SOTA video edit methods.
@@
 CLIP-Text Align CLIP-Consistency Alignment Consistency Preservation
-TAV [34] 0.2988 0.9086 19.06 2.64 4.25
-FateZero [22] 0.2738 0.9324 20.63 26.17 30.62
-ControlVideo [37] 0.2916 0.9301 8.48 5.30 3.70
-FLATTEN [5] 0.2931 0.9249 14.82 17.99 11.11
-TokenFlow [9] 0.2910 0.9561 12.67 20.11 24.40
+TAV (Wu et al. 2023) 0.2988 0.9086 19.06 2.64 4.25
+FateZero (Qi et al. 2023) 0.2738 0.9324 20.63 26.17 30.62
+ControlVideo (Zhang et al. 2023b) 0.2916 0.9301 8.48 5.30 3.70
+FLATTEN (Cong et al. 2023) 0.2931 0.9249 14.82 17.99 11.11
+TokenFlow (Geyer et al. 2023) 0.2910 0.9561 12.67 20.11 24.40
 ASSD (Ours) 0.2899 0.9113 24.34 27.79 25.93
-4.3 Quantitative Experiment
+Experiments Setting
+We collected over 30 text-video pairs from public datasets like DAVIS (Perazzi et al.
+2016), WebVid (Bain et al.
+2021), and YouTube (Abu-El-Haija et al.
+2016), extracting 8 or 16 evenly spaced frames from each video.
+We used ModelscopeT2V and Zeroscope as our base models.
+For optimizing the latent code of one video, we employed the SGD optimizer with a learning rate of 0.1 for 200 steps, taking approximately 3 to 4 minutes on a single V100 GPU.
+To evaluate the performance of our proposed method, we evaluated our method against five SOTA video editing methods: Tune-A-Video (TAV) (Wu et al.
+2023) fine-tunes the denoising U-Net on a single video to learn the structural integrity of the original video.
+FateZero (Qi et al.
+2023) achieves zero-shot video editing via DDIM Inversion and attention manipulation.
+FLATTEN (Cong et al.
+2023) uses an external optical flow detection model for accurate attention manipulation.
+TokenFlow (Geyer et al.
+2023) ensures consistency by modifying keyframes and propagating changes throughout the video.
+ControlVideo (CV) (Zhang et al.
+2023b) leverages depth and edge information from the original video to control editing results.
+Figure 4: Abalation of ASSD, Weighted Attention Fusion and Joint Guidance with Stable Diffusion.
+In most cases, using all the components we propose achieves the best balance between editing and minimizing contaminations.
+Qualitative Experiment
+Fig.
+3 presents the qualitative comparison results between our method and other SOTA baselines.
+Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and unedited regions, while allowing for geometric transformation of the edited objects.
+When an editing method allows for significant geometric transformations of the edited objects, it typically introduces challenges in preserving non-edited regions.
+Conversely, methods that struggle with geometric transformations of edited objects generally achieve better preservation of non-edited regions.
+Existing methods often face one of two issues.
+ControlVideo and Tune-A-Video allow for geometric transformations of the original video to better match the target prompt but fail to preserve the non-edited parts of the original video.
+FateZero, FLATTEN, and TokenFlow achieve better preservation of non-edited regions but struggle to support significant geometric transformations.
+Our method strikes a better balance between supporting geometric transformations and preserving non-edited regions.
+Quantitative Experiment
 Auto Metrics.
-We utilized the CLIP model [23] for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score.
+We utilized the CLIP model (Radford et al.
+2021) for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score.
 CLIP Text Alignment Score is defined as the mean cosine similarity between the CLIP image embeddings of the frames in the edited video and the CLIP text embedding of the target text prompt.
-This metric measures how well the edited video aligns with the target text prompt.
+This metric measures how well the frames in the edited video aligns with the target text prompt.
 CLIP Frame Consistency Score is defined as the mean cosine similarity between the CLIP image embeddings of frames within the video.
@@
 1.
+Although our method achieved only average results on both evaluation metrics, it is worth noting that there is still a lack of appropriate automatic evaluation metrics.
+The CLIP Frame Consistency Score focuses solely on the semantic consistency between different frames and fails to measure the continuity of motion and the consistency of objects between consecutive frames.
+Similarly, the CLIP Text Alignment Score only assesses the semantic alignment between the frames in the video and the target text prompt, and does not measure the preservation of non-edited parts.
+Figure 5: Visualization of One-Step Prediction of $\bm{Z}_{0}$ by Different Models at Large and Small Time Steps.
+$\bm{Z}_{0}$ is computed from the one-step predicted noise, and since visualizing $\bm{Z}_{0}$ is more meaningful, we choose to visualize $\bm{Z}_{0}$.
+Figure 6: The process of editing videos with our method.
+The images on the left depict the gradual transformation of a jeep into a bus, while the images on the right illustrate the gradual transformation of a jeep into a Lamborghini.
+Figure 7: Ablation study for the $\lambda$ hyperparameter in Weighted Attention Fusion. By adjusting lambda, we can balance between the stability of the editing and the extent of the editing.
 User Study.
-Due to the current lack of appropriate evaluation metrics in the field of video generation and editing, automatic assessments based on CLIP often do not align well with the perceptual quality of videos.
-Therefore, we primarily validate the effectiveness of our method through user studies.
+Since automatic assessments based on CLIP often do not align well with the perceptual quality of videos, we primarily validate the effectiveness of our method through user studies.
 We invited 40 volunteers to participate in our user study.
-The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment with the target text prompt.
-(2) Inter-frame consistency of the edited video.
+The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment between the edited video and the target text prompt.
+(2) Inter-frame consistency and the continuity of motion between adjacent frames of the edited video.
 (3) Preservation of the original video’s motion patterns and unedited regions.
@@
 1, our method achieves comparable or superior performance across all three evaluation metrics compared to other SOTA methods.
-Figure 5: Abalation of ASSD, Weighted Attention Fusion and Joint Guidance with Stable Diffusion.
-In most cases, using all the components we propose achieves the best balance between editing and minimizing contaminations.
-4.4 Ablation Study
-We conduct ablation studies below to validate the effect of each proposed component.
+Since FateZero typically only edits the color and texture of objects and struggles with geometric transformations, as seen in Fig.
+3 of the qualitative experiments, it effectively facilitates the preservation of non-edited parts and object IDs, resulting in the best performance in Preservation.
+Ablation Study
+We conduct ablation studies to validate the effect of each proposed component and the choice of hyperparameter.
 As shown in the first row of experimental results in Fig.
-5, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video.
-The results in the second row show that by using ASSD instead of DDS, and smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality.
+4, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video.
+The results in the second row show that by using ASSD to adaptively smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality.
 When we further incorporate the proposed Weighted Attention Fusion mechanism, using features from the reference branch, the contaminations are further reduced.
 Additionally, we observe that the edited video retains more information from the original video, such as colors and other details.
-Due to the significantly weaker robustness and generalization to different text conditions of current open-source text-to-video models compared to Stable Diffusion, the latter can sometimes provide higher-quality update gradients for certain text prompts, despite lacking the ability to model temporal information.
+Due to the significantly weaker robustness and generalization to different text conditions of current open-source T2V diffusion models compared to Stable Diffusion, we introduced Image-based Joint Guidance as the supervision.
 As shown in the last row of Fig.
-5, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects.
-5 Conclusion and Limitations
-We propose a new score distillation-based paradigm for video editing and introduce a novel score distillation method called Adaptive Sliding Score Distillation (ASSD).
-ASSD mitigates errors in the gradient information provided by text-to-video models through an adaptive sliding window-based smoothing operation and adaptively extracted masks.
-Our pipeline also incorporates auxiliary guidance from Stable Diffusion to provide more precise gradients for updating the latent code of original video.
-Additionally, to better capture the information from the original video to correct errors in the edited video and further improve the quality of the update gradients, we employ proposed weighted attention fusion.
-This modifies certain spatial self-attention layers in pre-trained text-to-video models to support the fusion of information between the editing branch and the reference branch.
-Experimental results validate the effectiveness and potential of our new paradigm in video editing tasks, as well as the efficacy of our proposed components.
+4, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects.
+In Fig.
+7, we validate different choices for the $\lambda$ hyperparameter in Weighted Attention Fusion.
+When $\lambda$ is large, more information from the reference branch is introduced into the editing branch, reducing contaminations in the output but also weakening the degree of editing.
+Conversely, when $\lambda$ is small, the opposite effect is observed.
+Conclusion and Limitations
+We introduce a novel score distillation-based paradigm for video editing, featuring Adaptive Sliding Score Distillation (ASSD).
+ASSD mitigates gradient errors from T2V diffusion models through adaptive sliding window-based smoothing and extracted masks.
+Our pipeline incorporates joint guidance from Stable Diffusion for precise latent code updates.
+We employ weighted attention fusion, modifying spatial self-attention layers in pre-trained T2V diffusion models to fuse information between editing and reference branches.
+Experimental results validate the efficacy of our paradigm and its components in video editing tasks.
 Limitations
-It should be noted that the performance of our method heavily relies on the capabilities of the text-to-video models providing the gradient information for editing.
-The effectiveness of such methods relies on the premise that the pre-trained text-to-video model can provide an update gradient that, despite containing errors, is still effective.
-Given the current lack of sufficiently powerful open-source text-to-video models, the upper bound of our method’s performance is significantly constrained.
+The efficacy of these methods hinges on the assumption that pre-trained T2V diffusion models provide effective update gradients, despite inherent errors.
+However, the dearth of robust open-source T2V diffusion models significantly constrains our method’s performance ceiling.